two planes, which are 1680 miles apart, fly toward each other. Their speed differ by 60mph. If they pass each other in 3hours, what is the speed of each?
Answer by macston(5194)

1680 miles/3hours=560 mph The combined speed of the two planes is 560 mph.
r=speed of slower plane; r+60mph=speed of faster plane
r+r+60mph=560mph
2r=500mph
r-250mph ANSWER 1: The speed of the slower plane is 250mph.
r+60mph=250mph+60mph=310mph ANSWER 2: The speed of the faster plane is 310 mph.